# From Restless to Contextual: A Thresholding Bandit Approach to Improve Finite-horizon Performance

### v2

@@ Appendix C Technical Details In Section 3
This will give us the desired result. Assuming Eq. (C.61) holds for $t=h$, we will select the greedy policy for $\mathcal{S}_{i}^{M-2}$ as we did in case 1. Again, the only case that is different from the 2-agent case is when $\bm{s}\in\mathcal{S}_{1}^{M-2},\mathcal{S}_{3}^{M-2}$. Therefore, we will only show Eq. (C.59) and Eq. (C.60) for those two cases when $t=h$ under the assumption that Eq. (C.59) and Eq. (C.60) hold for $t=h+1$ and Eq. (C.61) holds for $t=h$. We will first prove the RHS of Eq. (C.59). When $\bm{s}\in\mathcal{S}_{1}^{M-2}$, we have
@@
Now, we are going to prove the LHS of Eq. (C.59). When $\bm{s}\in\mathcal{S}_{1}^{M-2}$, we have
@@
Using the same calculation, we can show the case for $\bm{s}\in\mathcal{S}_{3}^{M-2}$ and Eq. (C.60).
@@
$\forall t\leq T,$ $0\leq V_{t}^{*}\left((1,0,\bm{s})\right)-V_{t}^{*}((0,0,\bm{s}))\leq{R(1)-R(0)}$ (C.62)
$0\leq V_{t}^{*}\left((1,1,\bm{s})\right)-V_{t}^{*}((1,0,\bm{s}))\leq R(1)-R(0)$ (C.63)
$V_{t}^{*}\left((1,1,\bm{s})\right)-V_{t}^{*}((1,0,\bm{s}))\leq V_{t}^{*}\left((1,0,\bm{s})\right)-V_{t}^{*}((0,0,\bm{s}))$ (C.64)
$\pi_{t+1}^{\text{greedy}}=\pi_{t+1}^{*}.$ (C.65)
When $t=T$, the statements hold trivially. Suppose this holds for $t\geq h+1$, when $t=h$, we will first show Eq. (C.65) under the assumption that Eq. (C.62)-(C.64) hold for $t=h+1$. Using the same reasoning, we only need to show Eq. (C.58). Again, from the above assumptions and the definition of $F$, we know that
00 0 $\leq F_{\bm{a}_{h+1},h+1}((1,1))-F_{\bm{a}_{h+1},h+1}((1,0))\leq{R(1)-R(0)}$
00 0 $\leq F_{\bm{a}_{h+1},h+1}((1,0))-F_{\bm{a}_{h+1},h+1}((0,0))\leq{R(1)-R(0)}.$
Then combined with the fact that ${{q^{1}_{11}q^{0}_{01}\geq q^{1}_{01}q^{0}_{11}}}$, it is easy to see that Eq. (C.58) holds.
Using the same logic used in the first sub-case, we can prove Eq. (C.62),(C.63) for $t=h$. The only thing we need to show is Eq. (C.64) for $t=h$. Again, when $\bm{s}\in\mathcal{S}_{2}^{M-2}$, this is the same as the two agent case. Therefore, we will only prove the other two cases under the assumption that Eq. (C.62)-(C.64) hold for $t=h+1$ and Eq. (C.65) holds for $t=h$. We will prove the result for $\bm{s}\in\mathcal{S}_{1}^{M-2}$, the other cases will hold for similar reasons. We have
@@
where the last inequality holds because of the assumption that Eq. (C.64) holds for $t=h+1$.
@@
$\forall t\leq T,$ $\frac{R(1)-R(0)}{q^{1}_{10}+q^{1}_{01}}\left(1-\left(q^{1}_{11}-q^{1}_{01}\right)^{T-t+1}\right)\leq V_{t}^{*}\left((1,0,\bm{s})\right)-V_{t}^{*}((0,0,\bm{s}))$ (C.66)
$\frac{R(1)-R(0)}{q^{0}_{10}+q^{0}_{01}}\left(1-\left(q^{0}_{11}-q^{0}_{01}\right)^{T-t+1}\right)\geq V_{t}^{*}\left((1,0,\bm{s})\right)-V_{t}^{*}((0,0,\bm{s}))$ (C.67)
$\frac{R(1)-R(0)}{q^{1}_{10}+q^{1}_{01}}\left(1-\left(q^{1}_{11}-q^{1}_{01}\right)^{T-t+1}\right)\leq V_{t}^{*}\left((1,0,\bm{s})\right)-V_{t}^{*}((0,0,\bm{s}))$ (C.68)
$\frac{R(1)-R(0)}{q^{0}_{10}+q^{0}_{01}}\left(1-\left(q^{0}_{11}-q^{0}_{01}\right)^{T-t+1}\right)\geq V_{t}^{*}\left((1,0,\bm{s})\right)-V_{t}^{*}((0,0,\bm{s}))$ (C.69)
$\pi_{t+1}^{\text{greedy}}=\pi_{t+1}^{*}.$ (C.70)
The proof will be exactly using the idea used in 2 agent case and the calculation for multiple agents case discussed above, thus will be omitted. ∎
Appendix D Technical Details In Section 4
We will continue to use the notation defined in Appendix C.
D.1 Proof of Lemma 4.3
@@ Appendix D Technical Details In Section 4
Before proving Theorem 4.4, we first introduce Lemma D.1 which decomposes the cumulative regret to per-timestep regret.
Lemma D.1.
@@
By Lemma D.1, we only need to bound $\text{Regret}_{t}(\pi,\mathcal{H}_{t})$. We have
@@
$\leq\mathbb{E}\left[\sum_{\Delta_{s}^{m}>0}\Delta_{s}^{m}\mathbbm{1}\left\{a_{t+1}^{m}=1,s_{t}^{m}=1\right\}\right]+\mathbb{E}_{\pi_{t+1}}\left[V_{t+1}^{\bm{\pi}^{b}}\left(\bm{s}_{t+1}\right)\mid\bm{s}_{t+1},\bm{a}_{t+1},\mathcal{H}_{t}\right]$
$=\mathbb{E}\left[\sum_{\Delta_{s}^{m}>0}\Delta_{s}^{m}\mathbbm{1}\left\{a_{t+1}^{m}=1,s_{t}^{m}=1\right\}\right],$
@@
$\sum_{\Delta_{s}^{m}>0}\Delta_{s}^{m}\mathbb{E}\left[\sum_{t=0}^{T}\mathbbm{1}\left\{a_{t+1}^{m}=1,s_{t}^{m}=s\right\}\right].$
@@
$\mathbb{E}\left[\sum_{t=0}^{T}\mathbbm{1}\left\{a_{t+1}^{m}=1,s_{t}^{m}=s\right\}\right]$ $=\mathbb{E}\left[\sum_{t=0}^{T}\mathbbm{1}\left\{a_{t+1}^{m}=1,s_{t}^{m}=s,\mathrm{LCB}_{t}^{m}(s)\geq\gamma\right\}\right]$ (D.71)
$+\mathbb{E}\left[\sum_{t=0}^{T}\mathbbm{1}\left\{a_{t+1}^{m}=1,s_{t}^{m}=s,\mathrm{LCB}_{t}^{m}(s)\leq\gamma\right\}\right].$ (D.72)
@@
We will bound the two terms separately. For Eq. (D.71), we have
$\mathbb{E}\left[\sum_{t=0}^{T}\mathbbm{1}\left\{a_{t+1}^{m}=1,s_{t}^{m}=s,\mathrm{LCB}_{t}^{m}(s)\geq\gamma\right\}\right]$
$=\mathbb{E}\left[\sum_{k=1}^{\infty}\mathbbm{1}\left\{\mathrm{LCB}_{\tau_{k}^{m}(s)-1}^{m}(s)\geq\gamma,\tau_{k}^{m}(s)\leq T\right\}\right]$
$\leq\mathbb{E}\left[\sum_{k=1}^{T}\mathbbm{1}\left\{\hat{I}_{\tau_{k}^{m}(s)-1}^{m}(s)\geq\gamma+\sqrt{\frac{3\log\left(N_{s,m}^{\left(\tau_{k}^{m}(s)\right)}\right)}{N_{s,m}^{\left(\tau_{k}^{m}(s)\right)}}}\right\}\right]$
$\leq\sum_{k=1}^{T}\exp\left(-N_{s,m}^{\left(\tau_{k}^{m}(s)\right)}\left(\Delta_{s}^{m}\right)^{2}+\log\log\left(N_{s,m}^{\left(\tau_{k}^{m}(s)\right)}\right)-3\log\left(N_{s,m}^{\left(\tau_{k}^{m}(s)\right)}\right)\right)$
$\leq\sum_{k=1}^{T}\frac{1}{k^{2}}\exp\left(-k\left(\Delta_{s}^{m}\right)^{2}\right)$
$\leq 2,$ (D.73)
where $\tau_{k}^{m}(s):=\inf\{t:N_{s,m}^{(t)}=k\}$ and the second inequality holds because of Lemma D.2.
For Eq. (D.72), we have
$\mathbb{E}\left[\sum_{t=0}^{T}\mathbbm{1}\left\{a_{t+1}^{m}=1,s_{t}^{m}=s,\mathrm{LCB}_{t}^{m}(s)\leq\gamma\right\}\right]$
@@
$=\mathbb{E}\left[\sum_{t=0}^{T}\mathbbm{1}\left\{a_{t+1}^{m}=1,s_{t}^{m}=s,\mathrm{LCB}_{t}^{m}(s)\leq\gamma,\exists(j,s_{t}^{j})\in\mathcal{G},\mathrm{LCB}_{t}^{j}(s_{t}^{j})\leq\gamma\right\}\right]$
$\leq\mathbb{E}\left[\sum_{t=0}^{T}\sum_{(j,s_{j})\in\mathcal{G}}\mathbbm{1}\left\{a_{t+1}^{m}=1,s_{t}^{m}=s,\mathrm{LCB}_{t}^{m}(s)\leq\gamma,\mathrm{LCB}_{t}^{j}(s_{j})\leq\gamma,s_{t}^{j}=s_{j}\right\}\right]$
$=\sum_{(j,s_{j})\in\mathcal{G}}\mathbb{E}\left[\sum_{t=0}^{T}\mathbbm{1}\left\{a_{t+1}^{m}=1,s_{t}^{m}=s,\mathrm{LCB}_{t}^{m}(s)\leq\gamma,\mathrm{LCB}_{t}^{j}(s_{j})\leq\gamma,s_{t}^{j}=s_{j}\right\}\right]$
For each $(j,s_{j})\in\mathcal{G}$, we have
@@
$\mathbb{E}\left[\sum_{t=0}^{T}\mathbbm{1}\left\{a_{t+1}^{m}=1,s_{t}^{m}=s,\mathrm{LCB}_{t}^{m}(s)\leq\gamma,\mathrm{LCB}_{t}^{j}(s_{j})\leq\gamma,s_{t}^{j}=s_{j}\right\}\right]$
@@
$=\mathbb{E}\left[\sum_{t=0}^{T}\mathbbm{1}\left\{a_{t+1}^{j}=1,s_{t}^{m}=s,\mathrm{LCB}_{t}^{m}(s)\leq\gamma,\mathrm{LCB}_{t}^{j}(s_{j})\leq\gamma,s_{t}^{j}=s_{j}\right\}\right]$
$\leq\mathbb{E}\left[\sum_{t=0}^{T}\mathbbm{1}\left\{a_{t+1}^{j}=1,\mathrm{LCB}_{t}^{j}(s_{j})\leq\gamma,s_{t}^{j}=s_{j}\right\}\right]$
@@
$=\mathbb{E}\left[\sum_{k=1}^{\infty}\mathbbm{1}\left\{\mathrm{LCB}_{\tau_{k}^{j}(s_{j})-1}^{m}(s_{j})\geq\gamma,\tau_{k}^{j}(s_{j})\leq T\right\}\right]$
$\leq\mathbb{E}\left[\sum_{k=1}^{T}\mathbbm{1}\left\{\hat{I}_{\tau_{k}^{m}(s_{j})-1}^{m}(s_{j})\geq\gamma+\sqrt{\frac{3\log\left(N_{s_{j},m}^{\left(\tau_{k}^{m}(s_{j})\right)}\right)}{N_{s_{j},m}^{\left(\tau_{k}^{m}(s_{j})\right)}}}\right\}\right]$
$\leq u+\mathbb{E}\left[\sum_{k=u+1}^{T}\mathbbm{1}\left\{\hat{I}_{\tau_{k}^{m}(s_{j})-1}^{m}(s_{j})\geq\gamma+\sqrt{\frac{3\log\left(N_{s_{j},m}^{\left(\tau_{k}^{m}(s_{j})\right)}\right)}{N_{s_{j},m}^{\left(\tau_{k}^{m}(s_{j})\right)}}}\right\}\right]$
$\leq u+\sum_{k=u+1}^{T}\frac{1}{k^{2}}$
$\leq u+2,$
where $u=\frac{12}{\left(\Delta_{s_{\min}}^{m_{\min}}\right)^{2}}\left(\log\left(\frac{e}{\left(\Delta_{s_{\min}}^{m_{\min}}\right)^{6}}\right)+\log\log\left(\frac{1}{\left(\Delta_{s_{\min}}^{m_{\min}}\right)^{6}}\right)\right)$. We should note that such choice of $u$ satisfies $\sqrt{3\log(u)/u}\leq\Delta_{s_{\min}}^{m_{\min}}$ (Lemma 18, Garivier & Kaufmann (2016)). Therefore, Eq. (D.72) can be upper bounded as
@@
$\mathbb{E}\left[\sum_{t=0}^{T}\mathbbm{1}\left\{a_{t+1}^{m}=1,s_{t}^{m}=s,\mathrm{LCB}_{t-1}^{m}(s)\leq\gamma\right\}\right]$
@@
$=\sum_{(j,s_{j})\in\mathcal{G}}\mathbb{E}\left[\sum_{t=0}^{T}\mathbbm{1}\left\{a_{t+1}^{m}=1,s_{t}^{m}=s,\mathrm{LCB}_{t-1}^{m}(s)\leq\gamma,\mathrm{LCB}_{t-1}^{j}(s_{j}),s_{t}^{j}=s_{j}\leq\gamma\right\}\right]$
$\leq|\mathcal{G}|(u+2).$ (D.74)
Combining Eq. (D.73) and Eq. (D.74) will give us the desired result.
Further assume that $k\left|\Delta_{s_{\min}}^{m_{\min}}\right|\geq\Delta_{s}^{m}$, for all $\Delta_{s}^{m}>0$. Then we have
@@
$\mathrm{R}(\pi)$ $\leq\sum_{\Delta_{s}^{m}\leq\Delta}\Delta_{s}^{m}\mathbb{E}\left[T_{s}^{m}\right]+\sum_{\Delta_{s}^{m}\geq\Delta}\frac{24k|\mathcal{G}|}{\Delta_{s_{\min}}^{m_{\min}}}\log\left(\frac{e}{\left(\Delta_{s_{\min}}^{m_{\min}}\right)^{6}}\right)+\sum_{\Delta_{s}^{m}>0}(2|\mathcal{G}|+2)\Delta_{s}^{m}$

### v4

@@ Appendix E Technical Details In Section 3
This will give us the desired result. Assuming Eq. (E.48) holds for $t=h$, we will select the greedy policy for $\mathcal{S}_{i}^{M-2}$ as we did in case 1. Again, the only case that is different from the 2-agent case is when $\bm{s}\in\mathcal{S}_{1}^{M-2},\mathcal{S}_{3}^{M-2}$. Therefore, we will only show Eq. (E.46) and Eq. (E.47) for those two cases when $t=h$ under the assumption that Eq. (E.46) and Eq. (E.47) hold for $t=h+1$ and Eq. (E.48) holds for $t=h$. We will first prove the RHS of Eq. (E.46). When $\bm{s}\in\mathcal{S}_{1}^{M-2}$, we have
@@
Now, we are going to prove the LHS of Eq. (E.46). When $\bm{s}\in\mathcal{S}_{1}^{M-2}$, we have
@@
Using the same calculation, we can show the case for $\bm{s}\in\mathcal{S}_{3}^{M-2}$ and Eq. (E.47).
@@
$\forall t\leq T,$ $0\leq V_{t}^{*}\left((1,0,\bm{s})\right)-V_{t}^{*}((0,0,\bm{s}))\leq{R(1)-R(0)}$ (E.49)
$0\leq V_{t}^{*}\left((1,1,\bm{s})\right)-V_{t}^{*}((1,0,\bm{s}))\leq R(1)-R(0)$ (E.50)
$V_{t}^{*}\left((1,1,\bm{s})\right)-V_{t}^{*}((1,0,\bm{s}))\leq V_{t}^{*}\left((1,0,\bm{s})\right)-V_{t}^{*}((0,0,\bm{s}))$ (E.51)
$\pi_{t+1}^{\text{greedy}}=\pi_{t+1}^{*}.$ (E.52)
When $t=T$, the statements hold trivially. Suppose this holds for $t\geq h+1$, when $t=h$, we will first show Eq. (E.52) under the assumption that Eq. (E.49)-(E.51) hold for $t=h+1$. Using the same reasoning, we only need to show Eq. (E.45). Again, from the above assumptions and the definition of $F$, we know that
$0$ $\leq F_{\bm{a}_{h+1},h+1}((1,1))-F_{\bm{a}_{h+1},h+1}((1,0))\leq{R(1)-R(0)}$
$0$ $\leq F_{\bm{a}_{h+1},h+1}((1,0))-F_{\bm{a}_{h+1},h+1}((0,0))\leq{R(1)-R(0)}.$
Then combined with the fact that $q^{1}_{11}q^{0}_{01}\geq q^{1}_{01}q^{0}_{11}$}}, it is easy to see that Eq. (E.45) holds.
Using the same logic used in the first sub-case, we can prove Eq. (E.49),(E.50) for $t=h$. The only thing we need to show is Eq. (E.51) for $t=h$. Again, when $\bm{s}\in\mathcal{S}_{2}^{M-2}$, this is the same as the two agent case. Therefore, we will only prove the other two cases under the assumption that Eq. (E.49)-(E.51) hold for $t=h+1$ and Eq. (E.52) holds for $t=h$. We will prove the result for $\bm{s}\in\mathcal{S}_{1}^{M-2}$, the other cases will hold for similar reasons. We have
@@
where the last inequality holds because of the assumption that Eq. (E.51) holds for $t=h+1$.
@@
$\forall t\leq T,$ $\frac{R(1)-R(0)}{q^{1}_{10}+q^{1}_{01}}\left(1-\left(q^{1}_{11}-q^{1}_{01}\right)^{T-t+1}\right)\leq V_{t}^{*}\left((1,0,\bm{s})\right)-V_{t}^{*}((0,0,\bm{s}))$ (E.53)
$\frac{R(1)-R(0)}{q^{0}_{10}+q^{0}_{01}}\left(1-\left(q^{0}_{11}-q^{0}_{01}\right)^{T-t+1}\right)\geq V_{t}^{*}\left((1,0,\bm{s})\right)-V_{t}^{*}((0,0,\bm{s}))$ (E.54)
$\frac{R(1)-R(0)}{q^{1}_{10}+q^{1}_{01}}\left(1-\left(q^{1}_{11}-q^{1}_{01}\right)^{T-t+1}\right)\leq V_{t}^{*}\left((1,0,\bm{s})\right)-V_{t}^{*}((0,0,\bm{s}))$ (E.55)
$\frac{R(1)-R(0)}{q^{0}_{10}+q^{0}_{01}}\left(1-\left(q^{0}_{11}-q^{0}_{01}\right)^{T-t+1}\right)\geq V_{t}^{*}\left((1,0,\bm{s})\right)-V_{t}^{*}((0,0,\bm{s}))$ (E.56)
$\pi_{t+1}^{\text{greedy}}=\pi_{t+1}^{*}.$ (E.57)
The proof will be exactly using the idea used in 2 agent case and the calculation for multiple agents case discussed above, thus will be omitted.
Case 4:
Suppose $q^{1}_{11}\leq q^{1}_{01},q^{0}_{11}\geq q^{0}_{01},q_{11}^{1}\leq q_{11}^{0}$. We can use the same way used in 2 agent case to directly calculate $2V_{t}^{\bm{\pi}^{\text{\text{greedy}}}}((1,0,\bm{s}))-V_{t}^{\bm{\pi}^{\text{\text{greedy}}}}((0,0,\bm{s}))-V_{t}^{\bm{\pi}^{\text{\text{greedy}}}}((1,1,\bm{s}))$ and have the result. ∎
Appendix F Technical Details In Section 4
We will continue to use the notation defined in Appendix E.
F.1 Proof of Lemma 4.3
@@ Appendix F Technical Details In Section 4
Before proving Theorem 4.4, we first introduce Lemma F.1 which decomposes the cumulative regret to per-timestep regret.
Lemma F.1.
@@
By Lemma F.1, we only need to bound $\text{Regret}_{t}(\pi,\mathcal{H}_{t})$. We have
@@
$\leq\mathbb{E}\left[\sum_{I^{m}(s)\leq\gamma}\Delta_{s}^{m}\mathbbm{1}\left\{a_{t+1}^{m}=1,s_{t}^{m}=1\right\}\right]+\mathbb{E}_{\pi_{t+1}}\left[V_{t+1}^{\bm{\pi}^{b}}\left(\bm{s}_{t+1}\right)\mid\bm{s}_{t+1},\bm{a}_{t+1},\mathcal{H}_{t}\right]$
$=\mathbb{E}\left[\sum_{I^{m}(s)\leq\gamma}\Delta_{s}^{m}\mathbbm{1}\left\{a_{t+1}^{m}=1,s_{t}^{m}=1\right\}\right],$
@@
$\sum_{I^{m}(s)\leq\gamma}\Delta_{s}^{m}\mathbb{E}\left[\sum_{t=0}^{T}\mathbbm{1}\left\{a_{t+1}^{m}=1,s_{t}^{m}=s\right\}\right].$
@@
$\mathbb{E}\left[\sum_{t=0}^{T}\mathbbm{1}\left\{a_{t+1}^{m}=1,s_{t}^{m}=s\right\}\right]$ $=\mathbb{E}\left[\sum_{t=0}^{T}\mathbbm{1}\left\{a_{t+1}^{m}=1,s_{t}^{m}=s,\mathrm{random}\right\}\right]$ (F.58)
$+\mathbb{E}\left[\sum_{t=0}^{T}\mathbbm{1}\left\{a_{t+1}^{m}=1,s_{t}^{m}=s,\mathrm{not~random}\right\}\right],$ (F.59)
@@
where $\mathrm{random}$ stands for doing random exploration. We will bound the two terms separately. For Eq. (F.58), we have
$\mathbb{E}\left[\sum_{t=0}^{T}\mathbbm{1}\left\{a_{t+1}^{m}=1,s_{t}^{m}=s,\mathrm{random}\right\}\right]$
@@
$=\mathbb{E}\left[\sum_{t=0}^{T}\mathbbm{1}\left\{a_{t+1}^{m}=1,s_{t}^{m}=s,\mathrm{random},\exists(m^{\prime},s^{\prime}),s_{t}^{m^{\prime}}=s^{\prime},N_{s^{\prime},m^{\prime}}^{(t)}\leq N_{s_{t}^{m^{\prime\prime}},m^{\prime\prime}}^{(t)},\forall m^{\prime\prime}\right\}\right]$
@@
$\leq\sum_{m^{\prime}}\sum_{s^{\prime}}\mathbb{E}\left[\sum_{t=0}^{T}\mathbbm{1}\left\{a_{t+1}^{m}=1,s_{t}^{m}=s,\mathrm{random},s_{t}^{m^{\prime}}=s^{\prime},N_{s^{\prime},m^{\prime}}^{(t)}\leq N_{s_{t}^{m^{\prime\prime}},m^{\prime\prime}}^{(t)},\forall m^{\prime\prime}\right\}\right]$
@@
$\leq\sum_{m^{\prime}}\sum_{s^{\prime}}\mathbb{E}\left[\sum_{t=0}^{T}\mathbbm{1}\left\{\mathrm{random},s_{t}^{m^{\prime}}=s^{\prime},N_{s^{\prime},m^{\prime}}^{(t)}\leq N_{s_{t}^{m^{\prime\prime}},m^{\prime\prime}}^{(t)},\forall m^{\prime\prime}\right\}\right]$
@@
$\leq\sum_{m^{\prime}}\sum_{s^{\prime}}\mathbb{E}\left[\sum_{k=0}^{\infty}\mathbbm{1}\left\{\mathrm{random},N_{s^{\prime},m^{\prime}}^{(\tau_{k}^{m^{\prime}}(s^{\prime}))}\leq N_{s_{t}^{m^{\prime\prime}},m^{\prime\prime}}^{(\tau_{k}^{m^{\prime}}(s^{\prime}))},\forall m^{\prime\prime},\tau_{k}^{m^{\prime}}(s^{\prime})\leq T\right\}\right]$
@@
$\leq\sum_{m^{\prime}}\sum_{s^{\prime}}\mathbb{E}\left[\sum_{k=0}^{T}\mathbbm{1}\left\{\mathrm{random},N_{s^{\prime},m^{\prime}}^{(\tau_{k}^{m^{\prime}}(s^{\prime}))}\leq N_{s_{t}^{m^{\prime\prime}},m^{\prime\prime}}^{(\tau_{k}^{m^{\prime}}(s^{\prime}))},\forall m^{\prime\prime}\right\}\right]$
$=\sum_{m^{\prime}}\sum_{s^{\prime}}\sum_{k=0}^{T}\mathbb{P}\left(\mathrm{random}\mid N_{s^{\prime},m^{\prime}}^{(\tau_{k}^{m^{\prime}}(s^{\prime}))}\leq N_{s_{t}^{m^{\prime\prime}},m^{\prime\prime}}^{(\tau_{k}^{m^{\prime}}(s^{\prime}))},\forall m^{\prime\prime}\right)\mathbb{P}\left(N_{s^{\prime},m^{\prime}}^{(\tau_{k}^{m^{\prime}}(s^{\prime}))}\leq N_{s_{t}^{m^{\prime\prime}},m^{\prime\prime}}^{(\tau_{k}^{m^{\prime}}(s^{\prime}))},\forall m^{\prime\prime}\right)$
$\leq\sum_{m^{\prime}}\sum_{s^{\prime}}\sum_{k=0}^{T}\mathbb{P}\left(\mathrm{random}\mid N_{s^{\prime},m^{\prime}}^{(\tau_{k}^{m^{\prime}}(s^{\prime}))}\leq N_{s_{t}^{m^{\prime\prime}},m^{\prime\prime}}^{(\tau_{k}^{m^{\prime}}(s^{\prime}))},\forall m^{\prime\prime}\right)$
$\leq\sum_{m^{\prime}}\sum_{s^{\prime}}\sum_{k=0}^{T}\frac{D}{k\Delta^{2}}$
$\leq SMD\frac{\log(T)}{\Delta^{2}},$ (F.60)
where $\tau_{k}^{m}(s):=\inf\{t:N_{s,m}^{(t)}=k\}$.
For Eq. (F.59), we have
$\mathbb{E}\left[\sum_{t=0}^{T}\mathbbm{1}\left\{a_{t+1}^{m}=1,s_{t}^{m}=s,\mathrm{not~random}\right\}\right]$
@@
$=\mathbb{E}\left[\sum_{t=0}^{T}\mathbbm{1}\left\{a_{t+1}^{m}=1,s_{t}^{m}=s,\mathrm{not~random},\exists(m^{\prime},s^{\prime})\in\mathcal{G},\hat{I}_{t}^{m}(s)\geq\hat{I}_{t}^{m^{\prime}}(s^{\prime}),s_{t}^{m^{\prime}}=s^{\prime}\right\}\right]$
$\leq\sum_{(m^{\prime},s^{\prime})\in\mathcal{G}}\mathbb{E}\left[\sum_{t=0}^{T}\mathbbm{1}\left\{a_{t+1}^{m}=1,s_{t}^{m}=s,\hat{I}_{t}^{m}(s)\geq\hat{I}_{t}^{m^{\prime}}(s^{\prime}),s_{t}^{m^{\prime}}=s^{\prime}\right\}\right]$
For each $(m^{\prime},s^{\prime})\in\mathcal{G}$, we have
$\mathbb{E}\left[\sum_{t=0}^{T}\mathbbm{1}\left\{a_{t+1}^{m}=1,s_{t}^{m}=s,\hat{I}_{t}^{m}(s)\geq\hat{I}_{t}^{m^{\prime}}(s^{\prime}),s_{t}^{m^{\prime}}=s^{\prime}\right\}\right]$
@@
$\leq\mathbb{E}\left[\sum_{t=0}^{T}\mathbbm{1}\left\{a_{t+1}^{m}=1,s_{t}^{m}=s,\hat{I}_{t}^{m}(s)\geq\gamma+\frac{\Delta}{2},s_{t}^{m^{\prime}}=s^{\prime}\right\}\right]$ (F.61)